Thursday, September 25, 2014

Station is permanently offline, replaced by CCMI2 buoy

The LCIY2 station as seen in 2011.

The CREWS station located just offshore from Little Cayman Research Centre in the Cayman Islands is now permanently offline.  It was destroyed while feeling the effects of wind and surf from Hurricane Sandy to the east as it crossed from Jamaica to Cuba on October 26, 2012.

This CREWS station was installed on July 22nd, 2009, through a collaboration between the Central Caribbean Marine Institute's (CCMI) Little Cayman Research Centre (LCRC) and the Atlantic Oceanographic and Meteorological Laboratory's (AOML) Coral Health and Monitoring Program (CHAMP).

The fiberglass tube is believed to have broken at its weakest point, where the internal polyurethane foam contained an air gap rather than providing continuous, uninterrupted reinforcement as per the station's design (see photo).

A closeup of the station's breaking point, showing where there was an air gap in the internally-reinforcing polyurethane foam.

During its lifetime this station was notable for surviving the effects of a magnitude 5.9 earthquake on January 19, 2010, whose epicenter was 45 miles ESE of Grand Cayman.  The station lost one of its lower support lines to the earthquake (see photo) but this was quickly repaired by LCRC personnel.

Minor damage from the earthquake on January 19, 2010.

Approximately one year after this station's loss, it was replaced by a new CREWS buoy at a location somewhat removed from this station's original site and in deeper water.  The new buoy is designated 42089 by NDBC and is known informally as CCMI2, as it is the second CREWS station to have been deployed by CCMI.

The CCMI2 station as seen shortly after its deployment in October of 2013.

There are not expected to be any further updates on this blog.  For updates on the CCMI2 station, please visit the new CCMI2 Field Log.

Thursday, October 11, 2012

Full Validation and Clean 9 October 2012

Cleaning of the CREWS station has been on schedule but the blog has fallen behind the last couple of months.  All dates have been blogged.  Apart from the cleaning, there is nothing really to report, we have had no storms, so the orientation, the lines, chains and equipment are all doing well.  Bio-fouling has been low over the summer, but we have had some fire coral creep in that we are tying to stay on top of.

The GT Sensor was installed at 1pm and cleaning commenced with two divers at 2:10pm.  The dive lasted about one and a half hours the instruments were cleaned first then the lines, chain and pole.  The GT sensor was removed at the end of the dive.  We also replaced the copper filters on the CTD.

Full Validation and Clean 14 September 2012

Full Validation and Clean 9 August 2012

Full Validation and Clean 11 July 2012

Friday, June 15, 2012

Full Validation and Clean 13 June 2012 and Nav light replacement

After the Ground Truth sensor was installed at 7:50, a cleaning session involving 7 divers started at 9:00.  All the instruments, chains, lines and pole were cleaned.  Bio-fouling was moderate, probably due to us cleaning a the station a couple of weeks late.  Cleaning finished at 11:30 and the GT Sensor removed at 2pm.

We tried to replace the anodes on the chain but between the two sizes we have, one is two big and another two small. We will source from somewhere else.  Three of the chains do now have new anodes.  Anodes and filters on the instruments were also cleaned and checked and are in good condition.  While we were out there we mounted a YSI pH instrument to the side of the pole one meter below the surface.  This will be collected in 2 weeks and the data used for one of CCMI's Ocean Acidification Manipulative studies.

A few weeks ago the Navigation Light on the Station stopped working, yesterday (June 14th) we replaced the light with a Carmanah M650, a newer model of the M601 which was installed before. ( please see photo)

Monday, May 07, 2012

Full Validation and Clean 2 May 2012

Due to an unexpected Nor'wester our cleaning of the CREWS station was postponed to and completed on 2nd May.  3 divers started cleaning at 9:15 and finished at 11:00. The GT sensor was installed at 7:50am and removed at 1:30.  Strong current made cleaning difficult near the surface but all was completed.  The sensors, lines, chain, and stick looked to be in good condition with low biofouling, no change on the orientation of the station.  It was noted that there was minimal fire coral present before cleaning.